# Retrieval Augmented Generation for Dynamic Graph Modeling

### v1

@@ 5.3 Model Analysis
To investigate the influence of the number of demonstrations, we conduct experiments across varying values $K\in\{1,3,5,7,9\}$. As shown in Fig. 5, a higher number of $K$ yields better prediction performance, that’s because more demonstrations provide richer contextual information, especially in the UCI dataset. However, including too many cases may introduce more noise, which can harm the performance.
Figure 5: Effect of the number of demonstrations K.
Effect of Different Fusion Strategies.
@@
To further investigate the effectiveness of the fusion strategy for the top-K demonstrations, we conduct experiments with different fusion strategies in Table 5 for K $=7$. We report NDCG@5 in Table 4 and present the other metrics in Appendix C. “Concatenation” denotes we directly concatenate the sequences of retrieved demonstrations and prepend them with the query sample sequence and then feed it into the pre-trained SimpleDyG model. “MLP” means we do not consider the graph structure of the demonstrations and replace the graph fusion as an MLP layer (we set the number of the MLP layer as 2). By using the MLP layer, We map the concatenated demonstrations into shorter m-dimensional embeddings (we empirically set m to be 15) and then concatenate it with the query sample. Like graph fusion, we only fine-tune the parameters of the MLP and output layer. The results in Table 5 show that directly concatenating the retrieved demonstrations with the query sample leads to lower performance compared with other strategies. This is because simple concatenation introduces a lengthy context, which can overwhelm the model with irrelevant information, and it neglects the structural relationships inherent in the demonstrations. The “MLP” strategy improves upon this by mapping the concatenated demonstrations into a shorter feature space, effectively reducing noise and emphasizing more relevant features. This approach yields better results than simple concatenation but still falls short compared to the “GraphFusion” strategy. The superior performance of the “GraphFusion” strategy highlights the importance of considering both the content and the structure of the demonstrations in the fusion process.
Fusion strategy UCI Hepth MMConv
Concatenation 0.036±0.018 0.007±0.002 0.110±0.030
MLP 0.106±0.017 0.015±0.002 0.167±0.006
GraphFusion 0.122±0.014 0.045±0.003 0.208±0.005
Table 5: Effect of different fusion strategies.
6 Conclusion
@@
In this work, we introduced a novel retrieval augmented framework RAG4DyG for dynamic graph modeling, addressing the limitations of existing approaches that often rely on narrow historical contexts. We leverage the RAG technique to broaden the context by incorporating relevant auxiliary information by harvesting high-quality demonstrations for dynamic graph samples and effectively utilizing this demonstration to enhance dynamic graph modeling. Our proposed solution involves a time-aware contrastive learning model to identify and retrieve temporally pertinent cases for each query sequence, coupled with a graph fusion strategy to integrate the inherent historical context with extended temporal contexts. Extensive experiments on real-world datasets across different domains demonstrated the superior performance of RAG4DyG in dynamic graph modeling.
Appendix
Appendix A Retrieval Data Annotation
To facilitate retrieval model training, we regard the samples in the training dataset as our retrieval pool $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ where each pair $(x_{i},y_{i})$ represents the historical sequence and its corresponding target sequence. Specifically, $x_{i}$ is the input sequence before the last time step and $y_{i}$ is the output sequence at the last time step. We annotate demonstrations based on the Jaccard similarity of $y_{i}$ among all the pairs in $D$.
$r(y_{i},y_{j})=\frac{|y_{i}\cap y_{j}|}{|y_{i}\cup y_{j}|}$ (11)
To control the quality of annotated data, we set a threshold of 0.8 to select highly similar demonstrations for each sample. These filtered annotations are then used to train the retriever model. The number of training samples for UCI, Hepth and MMConv datasets are 9,578, 8,250 and 10,762, respectively.
Appendix B More Details about Baselines
The details of the baselines are as follows:
  • DySAT (Sankar et al. 2020) utilizes self-attention mechanisms to capture both structural and temporal patterns in dynamic graphs through discrete-time snapshots.
  • EvolveGCN (Pareja et al. 2020) leverages recurrent neural networks to evolve the graph convolutional network parameters over discrete time steps.
  • DyRep (Trivedi et al. 2019) models dynamic graphs in continuous time by incorporating both temporal point processes and structural dynamics to capture interactions and node dynamics.
  • JODIE (Kumar, Zhang, and Leskovec 2019) focuses on user and item embedding trajectories over continuous time, predicting future interactions by modeling user and item embeddings jointly.
  • TGAT (Xu et al. 2020) employs temporal graph attention layers and time encoding to capture temporal dependencies and structural information for dynamic graphs.
  • TGN (Rossi et al. 2020) combines GNNs with memory modules to maintain node states over continuous time, effectively learning from dynamic interactions.
  • TREND (Wen and Fang 2022) integrates temporal dependencies based on Hawkes process and GNN to learn the dynamics of graphs.
  • GraphMixer (Cong et al. 2022) introduces a novel architecture that leverages MLP-mixer to learn link-encoder and node encoder for evolving graphs in continuous time.
  • SimpleDyG (Wu, Fang, and Liao 2024) reformulated the dynamic graph modeling as a sequence modeling task and mapped the dynamic interactions of target nodes as sequences with specially designed tokens. It simplifies dynamic graph modeling without complex architectural changes to effectively capture temporal dynamics.
Appendix C More Experimental Results
For the effect of different retrieval methods for generative performance in Sec. 5.3, we report the NDCG@5 score in Table 4. Here we report the remaining metrics in Table 6. The results show that RAG4DyG performs better than BM25 and Jaccard on Recall@5 and Jaccard metrics, which are consistent with the performance of NDCG@5.
Method UCI Hepth MMConv
Recall@5 Jaccard Recall@5 Jaccard Recall@5 Jaccard
BM25 0.111±0.007 0.093±0.004 - - 0.193±0.003 0.192±0.003
Jaccard 0.104±0.009 0.088±0.010 - - 0.193±0.004 0.192±0.004
RAG4DyG 0.111±0.013 0.097±0.010 0.019±0.002 0.019±0.002 0.194±0.005 0.194 ±0.005
GroundTruth 0.121±0.010 0.107±0.012 0.028±0.004 0.028±0.004 0.210±0.008 0.208±0.008
Table 6: Generative performance of retrieval methods.
For the effect of different fusion strategies in Sec. 5.3, the performance of the remaining metrics is shown in Table 7. Similarly, the “GraphFusion” strategy outperforms direct concatenation and “MLP”, indicating the effectiveness of the graph structure among demonstrations.
Fusion strategy UCI Hepth MMConv
Recall@5 Jaccard Recall@5 Jaccard Recall@5 Jaccard
Concatenation 0.033±0.019 0.029±0.016 0.001±0.002 0.002±0.002 0.103±0.031 0.103±0.031
MLP 0.102±0.018 0.089±0.016 0.006±0.002 0.006±0.002 0.156±0.006 0.153±0.006
GraphFusion 0.111±0.013 0.097±0.010 0.019±0.002 0.019±0.002 0.194±0.005 0.194 ±0.005
Table 7: Effect of different fusion strategies.
Appendix D Hyper-parameter Settings
We run all the experiments on an NVIDIA GPU L40 with the same data splitting and follow the hyper-parameter settings reported in SimpleDyG (Wu, Fang, and Liao 2024) for all the baselines. For our RAG4DyG method, we report the hyper-parameters as follows.
The time decay rate $\lambda$ for the retrieval model in Eq. (4) was fine-tuned according to the time granularity of different datasets, with days for the UCI dataset, months for the Hepth dataset, and turns for the MMConv dataset. We explored a range of values $\lambda=\{1e-4,1e-3,1e-2,1e-1,1,10,100\}$, ultimately selecting $\lambda=1e-4$ for UCI, $\lambda=0.1$ for Hepth, and $\lambda=10$ for MMConv. The coefficient $\alpha$ in the loss function (Eq. (8) was tuned across $\{0.2,0.4,0.6,0.8,1\}$, resulting in final values of $\alpha=1$ for UCI and MMConv dataset, $\alpha=0.4$ for Hepth. Additional parameter settings for the three datasets are as follows: the temperature $\tau$ in the two contrastive learning tasks is set to $\tau=\{0.1,0.1,0.1\}$; the batch size of the retriever model is $N=\{64,128,128\}$; and the masking and cropping portions in context-aware contrastive learning are $\{0.8,0.8,0.8\}$ and $\{0.4,0.6,0.6\}$, respectively.

### v2

@@ 5.3.3. Effect of the Number of Demonstrations K
To investigate the influence of the number of demonstrations, we conduct experiments across varying values $K\in\{1,3,5,7,9\}$. As shown in Fig. 5, a higher number of $K$ yields better prediction performance, that’s because more demonstrations provide richer contextual information, especially in the UCI dataset. However, including too many cases may introduce more noise, which can harm the performance. Ultimately, we select $K=7$ for all datasets.
Table 5. Effect of different fusion strategies.
Fusion strategy UCI Wikipedia Hepth Reddit
Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard
Concatenation 0.033±0.019 0.036±0.018 0.029±0.016 0.210±0.019 0.232±0.021 0.206±0.019 0.001±0.002 0.007±0.002 0.002±0.002 0.001±0.001 0.003±0.003 0.001±0.001
MLP 0.102±0.018 0.106±0.017 0.089±0.016 0.356±0.006 0.371±0.009 0.321±0.007 0.006±0.002 0.015±0.002 0.006±0.002 0.108±0.006 0.132±0.005 0.090±0.003
GraphFusion 0.111±0.013 0.122±0.014 0.097±0.010 0.369±0.006 0.389±0.008 0.328±0.007 0.019±0.002 0.045±0.003 0.019±0.002 0.119±0.006 0.143±0.005 0.096±0.003
5.3.4. Effect of Different Fusion Strategies.
@@
To further investigate the effectiveness of the fusion strategy for the top-K demonstrations, we conduct experiments with different fusion strategies under K $=7$. “Concatenation” denotes we directly concatenate the sequences of retrieved demonstrations and prepend them with the query sample sequence and then feed it into the pre-trained SimpleDyG model. “MLP” means we do not consider the graph structure of the demonstrations and replace the graph fusion as an MLP layer (we set the number of the MLP layer as 2). By using the MLP layer, We map the concatenated demonstrations into shorter m-dimensional embeddings (we empirically set m to be 15) and then concatenate it with the query sample. Like graph fusion, we only fine-tune the parameters of the MLP and output layer. The results in Table 5 show that directly concatenating the retrieved demonstrations with the query sample leads to lower performance compared with other strategies. This is because simple concatenation introduces a lengthy context, which can overwhelm the model with irrelevant information, and it neglects the structural relationships inherent in the demonstrations. The “MLP” strategy improves upon this by mapping the concatenated demonstrations into a shorter feature space, effectively reducing noise and emphasizing more relevant features. This approach yields better results than simple concatenation but still falls short compared to the “GraphFusion” strategy. The superior performance of the “GraphFusion” strategy highlights the importance of considering both the content and the structure of the demonstrations in the fusion process.
5.3.5. Time Complexity Analysis
The time complexity of our model RAG4DyG aligns with that of the vanilla Transformer, scaling as $O(n^{2})$, where $n$ denotes the sequence length. We measured the training time per epoch across various approaches using the UCI dataset to assess its efficiency. The results shown in Table 6 indicate that our model achieves faster or comparable training cost compared to the baseline methods. In contrast, approaches that integrate temporal components (e.g., RNNs or self-attention mechanisms) with structural elements (e.g., GNNs or GATs) face significant computational challenges due to the complexity of combining these modules.
Figure 6. Time efficiency and Recall@5 of different methods.
6. Conclusion
@@
In this work, we proposed RAG4DyG, a novel retrieval-augmented framework for dynamic graph modeling that overcomes the limitations of existing approaches by integrating broader temporal and contextual information. By leveraging the retrieval-augmented generation paradigm, RAG4DyG retrieves high-quality demonstrations from a retrieval pool and incorporates them effectively into the modeling process. The framework includes a time-aware contrastive learning module to prioritize temporally relevant samples and a graph fusion strategy to seamlessly integrate these retrieved demonstrations with the query sequence, enriching the historical context with extended temporal insights. Extensive experiments on diverse real-world datasets demonstrate the effectiveness of RAG4DyG in achieving state-of-the-art performance for dynamic graph modeling in both transductive and inductive scenarios.
Acknowledgements.
This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041 and T2EP20123-0052). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore. The authors would also like to thank HOANG Thi Linh for her assistance with this work.